OpenStudy (anonymous):

Suppose you deposit $5,000 in a savings account that earns 3% annual interest. If you make no other withdrawals or deposits, how many years will it take the account balance to reach at least $6,000? 10 years 6 years 7 years 4 years

OpenStudy (anonymous):

5000(1.03)^y >= 6000 Can you do logs?

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

ok, (1.03)^y >= 6/5 -> ln (1.03)^y >= ln 6/5 (y)(ln 1.03) >= ln 6/5 -> y >= (ln 6/5) / ln 1.03 y >= 6.168 So, with whole years, you need to get to 7 years.

OpenStudy (anonymous):

To verify this, you can take the equation in my first post with y=6 and with y=7. With y=6, you will see that you are short, but with y=7, you go over 6000.

OpenStudy (anonymous):

All good now?

OpenStudy (anonymous):

thanks!